OpenStudy (anonymous):

At a community picnic, the following liquids were combined to make a punch: 3 gallons of water 1.5 gallons of cranberry juice 1 pint of strawberry juice 1 cup of grape juice 3 quarts of pomegranate juice How much punch was made in all for the community picnic? Give your answer in cups. Select the best answer from the choices provided. 87 43.5 174 5.25

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (mrnood):

time you guys stopped using archaic measurement units and moved onto litres

OpenStudy (mrnood):

just type into google "3.5 galls in cups" anddo the same for the rest of the conversions - then add up the lot into 'cups'

OpenStudy (anonymous):

@MrNood I got 56

OpenStudy (mrnood):

that's correct for the water - now do the same for all the other ingredients convert to 'cups' and then add them up

OpenStudy (mrnood):

OOOPS soory - correction: It's 3 gallons water - not 3.5....

OpenStudy (zpupster):

add gallons first 3 + 1.5 + .75--(3 quarts are 3/4 gal. or .75) = xgal \[\frac{ xgal. }{ 1 }\left( \frac{ 16 cups }{ 1 gal} \right)\] the gallon units cancel now add 2cups because 2cups = pint add 1 more cup and that is your answer